OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = 2/7 and point (4, -1). A. y + 1 = -(2/7)(x – 4) B. y + 1 = (2/7)(x – 4) C. y – 1 = (2/7)(x – 4) D. y – 1 = -(2/7)(x – 4)

OpenStudy (texaschic101):

y - y1 = m(x - x1) y - (-1) = 2/7(x - 4) y + 1 = 2/7(x - 4)